# Synergizing Foundation Models and Federated Learning: A Survey

### v1

@@ 5.4 FM-FL for Healthcare
FMs, especially LLMs, have been found to excel in healthcare applications, showcasing impressive capabilities in tasks like mental health analysis Yang et al. (2023b), disease diagnosis Panagoulias et al. (2024), and drug discovery Chenthamarakshan et al. (2023). However, it raises privacy concerns to upload the health information of patients Tang et al. (2023) into a commercial server that supports the FMs. Meanwhile, FL has consistently received widespread attention in the healthcare domain Lincy and Kowshalya (2020); Rieke et al. (2020); Joshi et al. (2022), driven by the need for collaborative model training across different medical institutions without compromising patient data privacy. By breaking the barriers of private data availability, the FM-FL paradigm shows the potential to further harness the power of FMs in the healthcare domain.
@@
A recent study Shin et al. (2023a) presents a mobile mental health monitoring system, FedTherapist, which leverages user speech and keyboard input to fine-tune FMs with FL, demonstrating superior accuracy in mental health prediction tasks such as depression, stress, and mood prediction. Another representative study Feng et al. (2023a) focuses on Magnetic Resonance Imaging (MRI) reconstruction, which involves retrieving a complex-valued image from its under-sampled signal. The authors adopted an FM pre-trained on public datasets and trained visual prompts from decentralized clinical datasets via a personalized FL mechanism, thereby reducing communication costs and achieving competitive performance on limited local data.
@@
Despite the efforts, it has been shown that FMs in healthcare risk generating misleading information due to their imperfect understanding of complex medical data Jeblick et al. (2024).
6 Future Directions
@@
Multimodal FM-FL
With the development of mobile technology and IoT infrastructures Brunete et al. (2021), numerous edge devices produce data from a range of modalities, such as sensory, visual, and audio. In the era of FMs, the success of LLMs and their multimodal derivatives Ramesh et al. (2021); Google (2023); OpenAI (2024) have demonstrated the potential of multimodal FMs. The potential opportunities and challenges for multimodal FM-FL have yet to be explored.
@@
Continual learning enables models to adapt to new data over time, improving their performance and accuracy. By incorporating new data into the model training process, FL and FMs can continuously improve and adapt to changing environments and user needs Yang et al. (2024a). Future directions may involve leveraging transfer learning techniques in continual learning for FL and FMs. Models can transfer knowledge from previous tasks or domains to new ones, enabling more efficient adaptation Good et al. (2023).
@@
In scenarios where gradient access is unavailable, preliminary efforts have focused on federated fine-tuning black-box FMs Lin et al. (2023); Sun et al. (2024a); Lu et al. (2023b); Rui et al. (2024) utilizing ZOO. However, ZOO’s noticeably slower convergence rates, especially in high-dimensional contexts compared to gradient-based methods Golovin et al. (2020), indicate an important direction for further research. The impact of these slower convergence rates on overall efficiency and computational load within FL, particularly concerning large-scale FMs, has not been adequately investigated and understood.
@@
AI-Generated Content (AIGC) denotes content produced via advanced generative FMs Wu et al. (2023a). The strong generative capability of FMs offers the advantage of rapidly automating the creation of inexhaustible synthetic data. This capability positions AIGC as a valuable supplementary data source for model training and evaluation in many tasks Xu et al. (2024c). Despite some efforts Zhang et al. (2023b), more potential opportunities and challenges for AIGC-aided FL have yet to be explored.
7 Conclusions
@@
Limitations
FM and FL are very fast-moving fields. We have put a lot of effort into including the latest research efforts in the community in this survey. Therefore, we believe that our survey will help to inspire and push further research and innovation in these important areas. Our survey does not focus on experimental evaluation of the available ideas and systems. We believe that would be an important next step that we are leaving for future work.
Appendix A Additional Details of Adapter Tuning
A.1 Adapter Tuning
Adapter tuning integrates small-scale neural networks (known as “adapters”) into the pre-trained models Houlsby et al. (2019); Hu et al. (2022). A straightforward implementation of adapter tuning is to collaboratively train a shared adapter among all clients in the FedAvg manner, as highlighted by Sun et al. (2022a). Based on FedAvg, FedCLIP Lu et al. (2023a) incorporates an attention-based adapter for the image encoder in CLIP models Radford et al. (2021). In the domain of multilingual machine translation, where different language pairs exhibit substantial discrepancies in data distributions, Fed-MNMT Liu et al. (2023d) explores clustering strategies that group adapter parameters and makes inner-cluster parameters aggregation for alleviating the undesirable effect of data discrepancy. Another representative approach named C2A Kim et al. (2023) employs hypernetworks Ha et al. (2017) to generate client-specific adapters by conditioning on the client’s information, maximizing the utility of shared model parameters while minimizing the divergence caused by data heterogeneity.
A.2 Prompt Tuning
Prompt tuning incorporates trainable task-specific continuous prompt vectors at the input layer Liu et al. (2023a); Dong et al. (2023). Compared to full fine-tuning, it achieves comparable performance but with 1000×1000$\times$1000 × less parameter storage and communication Jia et al. (2022). A variation of prompt tuning, FedPerfix Sun et al. (2023) uses a local adapter to generate the prefixes and aggregate the original self-attention layers.
Depending on target modalities, prompt tuning in current literature can be further classified into three categories:
  • Textual Prompt Tuning. Task-specific prompt embeddings are combined with the input text embeddings, which are subsequently fed into language models. These soft prompts serve as instructive contexts to influence the generation process of LLMs by steering the probability distribution of the next token Dong et al. (2023).
  • Visual Prompt Tuning. Taking inspiration from advances in efficiently tuning LLMs, prompts are also introduced in the input space of vision models Jia et al. (2022). Naive implementations introduce prompts at the pixel level, acting as a form of data augmentation Li et al. (2024a). Alternatively, one could also insert the prompts as latent vectors for the first Transformer layer Deng et al. (2024); Yang et al. (2023a). Nevertheless, an empirical study Jia et al. (2022) has suggested that it is easier for visual prompts to learn condensed task-dependent signals in the latent input space of Transformers.
  • Textual-Visual Prompt Tuning. Unlike single-modal FMs, vision-language FMs can process and interpret both visual data and textual information, endowing them with powerful representation ability and transferability Radford et al. (2021). Based on vision-language FMs like CLIP, textual-visual prompt tuning shows promising capabilities in FL Guo et al. (2023), especially in cross-domain scenarios, where the model needs to generalize across varied domains and unseen classes Qiu et al. (2024).
Table 3: A list of existing FM-FL libraries and benchmarks. Missing or inapplicable details denoted by N/A. ✓ denotes a strong focus or presence; ✗ indicates no focus or absence; ◐ signifies a moderate focus or partial inclusion.
Library/Benchmark FL Backend LLM Support MultiModal FM Support FedPEFT On-Device Training Distributed & Clustered Differential Privacy Description FederatedScope-LLM Kuang et al. (2023) FederatedScope An end-to-end benchmark for efficient fine-tuning LLMs with FL NVIDIA FLARE Roth et al. (2024) NVFlare Scalable and efficient fine-tuning LLMs with FL FATE-LLM Fan et al. (2023) FATE Focuses on IP and privacy protection in federated LLM FedLLM FedML (2023) FedML An MLOps-supported training pipeline based on FedML OpenFedLLM Ye et al. (2024) N/A N/A An LLM framework focusing on FL instruction tuning/alignment Shepherd Zhang et al. (2024b) N/A Federated instruction tuning based on Hugging Face FedPETuning Zhang et al. (2023f) FedLab A benchmark comprising four FedPEFT methods FedLegal Zhang et al. (2023e) FedLab A benchmark comprising six legal NLP tasks under FL settings
Appendix B Libraries and Benchmarks
This part briefly introduces a series of available libraries and benchmarks for developing and examining FM-FL techniques. An overview is provided in Table 3.
  • FederatedScope-LLM Kuang et al. (2023) is an open-source package for fine-tuning LLMs via FL. Built on top of a popular FL backend FederatedScope Xie et al. (2023), it supports federated fine-tuning of LLMs under various FL scenarios, including FedPEFT and model personalization.
  • NVIDIA FLARE Roth et al. (2024) is an FL framework that allows researchers and data scientists to seamlessly move their machine learning and deep learning workflows into a federated paradigm.
  • FATE-LLM Fan et al. (2023) is an industrial-grade FL framework for LLM. Apart from FedPEFT, it provides a privacy hub integrating several IP protection and privacy-preserving mechanisms to protect model security and data privacy.
  • FedLLM FedML (2023) is an MLOps-supported training pipeline built upon the FedML AI platform He et al. (2020). FedLLM is compatible with popular LLM libraries such as HuggingFace and DeepSpeed to support a large range of FMs and datasets.
  • OpenFedLLM Ye et al. (2024) is a federated tuning framework for LLMs, which covers applications of instruction tuning and value alignment, diverse FL baselines, training datasets, and evaluation datasets.
  • Shepherd Zhang et al. (2024b) is a lightweight federated tuning framework. The local training process of Shepherd is built upon the implementations of Alpaca-LoRA Wang (2023), and Hugging Face’s PEFT Mangrulkar et al. (2022), enabling efficient fine-tuning.
  • FedPETuning Zhang et al. (2023f) is a pioneering federated benchmark for four representative FedPEFT methods, covering adapter tuning, prefix tuning, LoRA, and BitFit.
  • FedLegal Zhang et al. (2023e) is the very first real-world FL benchmark for legal NLP, which comprises five legal NLP tasks and one privacy task based on the data from Chinese courts.

### v2

@@ 8.4. FedFM for Healthcare
FMs, especially LLMs, have been found to excel in healthcare applications, showcasing impressive capabilities in tasks like mental health analysis (Yang et al., 2023b), disease diagnosis (Panagoulias et al., 2024), and drug discovery (Chenthamarakshan et al., 2023). However, it raises privacy concerns to upload the health information of patients (Tang et al., 2023) into a commercial server that supports the FMs. Meanwhile, FL has consistently received widespread attention in the healthcare domain (Lincy and Kowshalya, 2020; Rieke et al., 2020; Joshi et al., 2022), driven by the need for collaborative model training across different medical institutions without compromising patient data privacy. By breaking the barriers of private data availability, the FedFM paradigm shows the potential to further harness the power of FMs in the healthcare domain.
@@
A recent study (Shin et al., 2023a) presents a mobile mental health monitoring system, FedTherapist, which leverages user speech and keyboard input to fine-tune FMs with FL, demonstrating superior accuracy in mental health prediction tasks such as depression, stress, and mood prediction. Another representative study (Feng et al., 2023) focuses on Magnetic Resonance Imaging (MRI) reconstruction, which involves retrieving a complex-valued image from its under-sampled signal. The authors adopted an FM pre-trained on public datasets and trained visual prompts from decentralized clinical datasets via a personalized FL mechanism, thereby reducing communication costs and achieving competitive performance on limited local data.
@@
Despite the efforts, it has been shown that FMs in healthcare risk generating misleading information due to their imperfect understanding of complex medical data (Jeblick et al., 2024).
9. Future Directions
@@
Multimodal FedFM
With the development of mobile technology and IoT infrastructures (Brunete et al., 2021), numerous edge devices produce data from a range of modalities, such as sensory, visual, and audio. In the era of FMs, the success of LLMs and their multimodal derivatives (Ramesh et al., 2021; Google, 2023; OpenAI, 2024) have demonstrated the potential of multimodal FMs. The potential opportunities and challenges for multimodal FedFM have yet to be explored.
@@
Continual learning enables models to adapt to new data over time, improving their performance and accuracy. By incorporating new data into the model training process, FL and FMs can continuously improve and adapt to changing environments and user needs (Yang et al., 2024a). Future directions may involve leveraging transfer learning techniques in continual learning for FL and FMs. Models can transfer knowledge from previous tasks or domains to new ones, enabling more efficient adaptation (Good et al., 2023).
@@
In scenarios where gradient access is unavailable, preliminary efforts have focused on federated fine-tuning black-box FMs (Lin et al., 2023; Sun et al., 2024a; Lu et al., 2023b; Rui et al., 2024) utilizing ZOO. However, ZOO’s noticeably slower convergence rates, especially in high-dimensional contexts compared to gradient-based methods (Golovin et al., 2020), indicate an important direction for further research. The impact of these slower convergence rates on overall efficiency and computational load within FL, particularly concerning large-scale FMs, has not been adequately investigated and understood.
@@
AI-Generated Content (AIGC) denotes content produced via advanced generative FMs (Wu et al., 2023a). The strong generative capability of FMs offers the advantage of rapidly automating the creation of inexhaustible synthetic data. This capability positions AIGC as a valuable supplementary data source for model training and evaluation in many tasks (Xu et al., 2024c). Despite some efforts (Zhang et al., 2023b), more potential opportunities and challenges for AIGC-aided FL have yet to be explored.
10. Conclusions